# User Perception of Partially Automated Driving Systems: A Meaningful Human Control Perspective on the Perception among Tesla Users

### v1

@@ II-A Meaningful human control of automated driving systems
[5] provides a comprehensive philosophical account of two conditions for a system to be under meaningful human control. First, the tracking condition requires that a system should be capable of responding to relevant moral, strategic, and intentional reasons of humans in the environment where the system operates. Second, the tracing condition implies that a system should be designed in a way that allows tracing back the outcome of its operation to at least one human in the loop of control. [12] operationalized the tracking condition for partially automated driving systems, connecting strategic and tactical reasons with Michon’s classical theory of the driving task [13]. In this approach, the tracking condition is satisfied when the behavior of the automated driving system aligns with the moral (e.g. respecting regulations), strategic (e.g., going home), tactical (e.g., overtaking), and operational (e.g., steering) reasons of relevant humans (e.g., the driver). For instance, if a driver has a tactical reason to change lanes smoothly, the system should perform a lane change considered smooth by the driver. Findings from [12] have then been used as a basis for design guidelines on human-automation interaction in mixed-traffic [7]. Furthermore, [14] applied these findings to create a quantitative formulation for vehicle control.
@@ II-B User perception of safety and trust in automated driving systems
Studies consistently demonstrate that drivers’ perceptions, particularly their perceptions of safety and trust, significantly influence their willingness to adopt automated driving systems [15, 16]. However, such studies have so far been mostly limited to driving simulator experiments and public surveys [17, 18, 19], which warrants caution in assuming that the results accurately reflect real-world driving experiences. A systematic review of driving simulator experiments revealed variations in the representations [20]. Furthermore, many survey studies primarily relied on drivers’ imaginative perceptions and expectations of automated vehicles, lacking practical experience and detailed knowledge about the vehicles’ functionality [21]. Survey studies also have disadvantages in terms of information access, reliability, and validity [22]. For this reason, in this work we focus on semi-structured interview data collected previously [11] that avoids these issues but also provides in-depth insight into the tactical reasons of users.
@@ III-B Data
In this study, we utilized transcribed conversation data from [11]. Our analysis covered aspects that were not discussed in previous research that used this data [11, 25, 23]. The data comprised responses from 103 respondents in the FSD Beta program, collected through interviews using a semi-structured protocol that included both open-ended and closed-ended questions. The questions consisted of five sections: general experience, perceptions of vehicle operation, perceptions of safety, exploration of trust level, and typical vehicle usage. The interviews were conducted via Zoom where the participants were recruited through social media. Following quality checks, data from only 99 participants were deemed suitable for further analysis in this research. We only used the answers to the open-ended questions because the purpose was to look for reasons that were only mentioned when the participants explained them. Throughout the interviews, both video and conversation were recorded. The interviewer’s role was limited to reduce the possibility of bias.
@@ III-C Data processing
Based on the collection of documents where the transcribed conversation data was saved, we further processed the data so that it could be used for further analysis in our keyword search algorithm. We processed each transcript by applying a series of text preprocessing steps: removing newline characters, adding spaces between digits and alphabets, and combining these operations to create a cleaned version of the conversation. The text was then tokenized using the ‘$word\_tokenize$‘ function in NLTK to break it into individual words. Subsequently, a text cleaning function was applied, removing short words, eliminating spurious characters, transforming letters to lowercase, and removing stopwords and specific words defined in the process, such as the names of participants. The final step involved lemmatization, which normalizes the tokens using NLTK’s WordNet lemmatizer.
To illustrate, consider a sample text such as ”So changing lanes and avoiding blind spot collisions is very, very good here. I think I’m maintaining speeds and safe speeds are very good.”. After applying the data processing, the lemmatized tokenized data results will be formatted as a list: [’changing’, ’lane’, ’avoiding’, ’blind’, ’spot’, ’collision’, ’good’, ’think’, ’maintaining’, ’speed’, ’safe’, ’speed’, ’good’].
@@ III-D Seed words and keyword search
To identify instances in the transcribed conversation data where interview participants mentioned the driving tasks and their perceived safety and trust, we employed seed words and a keyword search algorithm with the data. Seed words were defined as individual terms associated with driving tasks, perceived safety, and trust. For driving tasks, we defined the seed words as the verb corresponding to each driving task, except for steering. The choice of the seed word ”steer” was often connected to the expression ”steering wheel,” which wasn’t our focus. Therefore, we sought alternative terms describing the steering process, namely ’change lane’ and ’keep lane’. Defining seed words is an iterative step, aligning with [26]’s suggestion to derive seeds based on the researcher’s subject knowledge. For perceived safety and trust, we adapted the seed words used by [23]. Subsequently, our keyword search algorithm, which operates by using the defined seed words as keywords, was employed to identify instances in each conversation where participants mentioned these seed words. The seed words and the algorithm can be found in Algorithm 1 and Table I. In Table I, the use of ’and’ indicates that both words must be present for the keyword search algorithm, while the use of ’or’ indicates that the presence of either one is sufficient.
As an example, assume we have a list of lemmatized tokenized data: [’vehicle’, ’good’, ’brake’, ’really’, ’hard’, ’try’, ’make’, ’left’, ’turn’, ’reason’, ’turn’, ’blinker’, ’quick’, ’definitely’, ’make’, ’feel’, ’unsafe’, ’house’, ’spot’, ’signal’, ’make’, ’better’, ’before’]. We are looking for words discussing perceived safety and trust, particularly those related to braking. We set ’brake’ as the seed word for braking and ’safe,’ ’happy,’ ’relax,’ ’comfort,’ ’glad,’ ’trust,’ ’distract,’, ’rely,’ and ’trustworthy’ as seed words for perceived safety and trust. The algorithm goes through the list, searching for occurrences of these seed words. When it finds one occurrence in one of the seed words, for instance, the word ’brake’, it will expand the search over the next 20 words, resulting in a new list of words [’brake’, ’really’, ’hard’, ’try’, ’make’, ’left’, ’turn’, ’reason’, ’turn’, ’blinker’, ’quick’, ’definitely’, ’make’, ’feel’, ’unsafe’, ’house’, ’spot’, ’signal’, ’make’]. Then, the algorithm checks for any words related to the seeds of perceived safety and trust, and it identifies the word ’safe’ inside the term ’unsafe’. Now, there are two occurrences of seed words in the list, and the list will be retrieved.
Finally, we qualitatively classified perceptions by examining keywords and their location in transcribed conversations. This included checking whether the systems tracked the participant’s reasons during actions such as braking and determining if the braking process indicated a positive or negative perception of safety and trust.
@@ IV-A1 Controversy: lane changing
The participants mentioned thirteen instances where they felt safe with the systems. Nine instances indicated that the systems tracked participants’ reasons, while the rest did not. One participant emphasized that significant improvements in automatic lane changes contributed to their sense of safety. Previously, their vehicle would perform several unsafe lane changes, but now it has become as safe as their own driving. Another participant simply didn’t want to be bothered with lane-changing, as they believed that the systems worked properly and safely.
  • ”I do feel safe when it’s doing automatic lane changes, and that’s a massive improvement since when it first was released, it used like almost every autopilot lane change when it first came out. Would be cutting someone off. Pulling into a lane with someone rapidly approaching or some other form of unsafe lane change now. Uhm.The automatic lane change feature. Is like. I would say just about as safe as I am. (R007)”
@@
  • ”I feel safe when I am on the freeway going straight.. I feel very safe. I mean, it’s s better driver than I am at taking turns on the freeway.. Changing lanes.. I like how it keeps his turn signal on all the way through the entire lane change and sometimes human drivers will just do a couple blinks. (R043)”
@@ IV-A3 Reasons mostly tracked: lane keeping
When the participants perceived lane keeping as safe, all instances showed that their reasons were tracked. Participants mentioned eleven instances where the automated driving system’s lane keeping tracked their reasons and gave them safe perceptions. One participant described that they felt safe because the systems reduced their main workload so that they didn’t have to pay attention to lane keeping. Another participant mentioned that the systems did a very good job of lane keeping. For example:
  • ”I feel like Autopilot makes me safer because it reduces my workload.. in terms of Lane keeping mainly.. It makes me feel more comfortable as a driver just because I don’t have to pay attention to the lines on the road. It’s more like seeing threats down the road. (R024)”
@@ IV-B1 Reasons mostly not tracked: lane changing
When the participants explained the situations in which they felt unsafe while changing lanes, the automated driving systems did not track their reasons at all. One of the participants mentioned that they didn’t feel safe because the system did not have human-like behaviors. They further emphasized that it was too fast or too close to other vehicles at stop signs or traffic lights. Another participant highlighted that they could feel safer if the systems could adapt to how humans drive and incorporate a more human-like feeling into them.
@@ IV-D1 Reasons mostly not tracked: braking
Participants who felt lack of trust described that the systems failed to track their reasons. They mentioned that they did not trust the automated driving systems because the systems could make unexpected wrong decisions very fast, such as braking when there was a pedestrian on the crosswalks. Another participant described that they didn’t trust the systems when entering a highway because the systems didn’t brake when there was not much of a gap.
  • ”I don’t fully trust it when it’s active for the reasons I stated before where it can just do something wrong very fast.. you just do not expect it. For example, if it sees a pedestrian.. on the crosswalk, it’ll kind of just slam on the brakes. (R065)”
@@
  • ”Then I think the Autopilot may even be better than a human, and if we take emergency braking, then it can probably react and brake faster than me or certainly close to that..but as I said, you have to assess the situation, then it’s uncertain, and that’s why I always have to rely on myself. (R015)”
@@ V-A1 Perceived safety
On the other hand, feeling safe doesn’t always mean that the automated driving systems tracked all their reasons; there were instances of lane-changing and braking tasks where tracking failed. One participant noted that braking was too slow for them, indicating their expected deceleration reason wasn’t tracked. However, they felt safe as long as they were ready to take over. When the driver realized operational reasons for deceleration weren’t fulfilled, they often took control. The participant also reported no difficulties with the takeover process. [28] found that aggressive drivers were more likely to take control when driving defensively programmed automated vehicles (AVs), potentially explaining the frequent initiation of takeover processes by drivers. The study also revealed that perceived safety levels were similar when aggressive drivers operated both aggressive and defensive AVs. Despite this, the safety score remained higher than when defensive drivers operated aggressive AVs, potentially explaining the continued feeling of safety.
The other instances indicated that the vehicle acted differently from what they expected, but it still led to perceived safety because they believed it was a better driver. The experience of many unexpected maneuvers that turned out to be correct decisions also influenced their belief. The more the vehicles behaved according to the expected reasons, the more they gained trust from the driver. This trust would make people perceive the system as a safe, even if it fails to track their reasons. This is in line with the findings on the effect of reliable experiences with trust that the system is safe [27]. Thus, we found indications that tracking driver expectations, such as performing safe maneuvers like lane-keeping, braking, and lane-changing like human drivers, positively correlated with perceived safety. However, failure to track reasons for lane changing and braking wasn’t necessarily associated with perceived lack of safety; it depended on other factors like the numerous reliable experiences, the driver’s trust level, and ease of taking over control.
@@ V-B1 Trust
We found indications that tracking driver expectations, including improved maneuver execution and human-like performance in intended traffic situations in lane-changing and braking tasks, positively relates to trust. Additionally, transparency, which does not relate to tactical or operational reasons, was also positively associated with trust. Notably, the failure to track reasons for braking tasks didn’t always result in the lack of trust; factors such as positive experiences and the ease of taking over control played a crucial role here.
@@ V-C Limitations
First, the data in this research focused on drivers’ subjective perceptions which may not accurately reflect the actual on-road situation (e.g., findings from [35] indicate that subjective risk can significantly differ from actual crash risk). To address this, we recommend further research to investigate telemetry data in addition to subjective reports. Second, we limited the tactical and operational reasons for relevant human agents only to the driver’s perspective. However, many other agents (e.g., vehicle manufacturers and other road users) could and should have meaningful human control over automated driving systems [12]. We recommend further research to evaluate the tracking of the reasons of other relevant human agents for a more holistic analysis. Third, our results (e.g., Table II) might not have fully captured the entirety of the participants’ responses. The use of seed words may restrict instances that are essentially the same but articulated with different words. We recommend incorporating a more extensive set of alternative seed words, especially in the context of driving tasks, for more comprehensive results. Fourth, our findings are limited to the participants in the FSD Beta program. Results from [36] indicate differences in the number of activations of safety features and level of trust among different vehicles from various manufacturers. We recommend that future research include a broader representation of manufacturers and participants, ensuring a more diverse sample that better reflects the population.
@@ Acknowledgments
We declare that OpenAI ChatGPT was used in the writing process to enhance the style and conciseness of the text originally written by the authors. This work was funded by the Indonesia Endowment Fund for Education (LPDP) under Grant 0006552/TRA/D/19/lpdp2021.

### v2

# A Meaningful Human Control Perspective on User Perception of Partially Automated Driving Systems: A Case Study of Tesla Users

@@ II-A Meaningful human control of automated driving systems
Santoni de Sio & van den Hoven [5] provide a comprehensive philosophical account of two conditions for a system to be under meaningful human control. First, the tracking condition requires that a system should be capable of responding to relevant moral, strategic, and intentional reasons of humans in the environment where the system operates. Second, the tracing condition implies that a system should be designed in a way that allows tracing back the outcome of its operation to at least one human in the loop of control. The tracking condition has been operationalized for partially automated driving systems [12], connecting strategic and tactical reasons with Michon’s classical theory of the driving task [13]. In this approach, the tracking condition is satisfied when the behavior of the automated driving system aligns with the moral (e.g. respecting regulations), strategic (e.g., going home), tactical (e.g., overtaking), and operational (e.g., steering) reasons of relevant humans (e.g., the driver). For instance, if a driver has a tactical reason to change lanes smoothly, the system should perform a lane change considered smooth by the driver. Findings from [12] have then been used as a basis for design guidelines on human-automation interaction in mixed-traffic [7]. Furthermore, [14] applied these findings to create a quantitative formulation for vehicle control.
@@ II-B User perception of safety and trust in automated driving systems
Studies consistently demonstrate that drivers’ perceptions, particularly their perceptions of safety and trust, significantly influence their willingness to adopt automated driving systems [15, 16]. However, such studies have so far been mostly limited to driving simulator experiments and public surveys [17, 18], which warrants caution in assuming that the results accurately reflect real-world driving experiences. A systematic review of driving simulator experiments revealed variations in the representations [19]. Furthermore, many survey studies primarily relied on drivers’ imaginative perceptions and expectations of automated vehicles, lacking practical experience and detailed knowledge about the vehicles’ functionality [20]. Survey studies also have disadvantages in terms of information access, reliability, and validity [21]. For this reason, in this work we focus on semi-structured interview data collected previously [11] that avoids these issues but also provides in-depth insight into the tactical reasons of users.
@@ III-B Data
In this study, we utilized transcribed conversation data from [11]. Our analysis covered aspects that were not discussed in previous research that used this data [11, 24, 22]. The data comprised responses from 103 respondents in the FSD Beta program, collected through interviews using a semi-structured protocol that included a total of 35 questions, comprising both open-ended and closed-ended questions. The questions consisted of five sections: general experience, perceptions of vehicle operation, perceptions of safety, exploration of trust level, and typical vehicle usage. The interviews were conducted via Zoom and the participants were recruited through social media. On average, each interview lasted 78 minutes and yielded approximately 12,200 words. Following quality checks, which involved the removal of non-English interview data and addressing missing transcriptions, only 99 interview data were deemed suitable for further analysis in this research. We only used the answers to the open-ended questions because the purpose was to look for reasons that were only mentioned when the participants explained them. Throughout the interviews, both video and conversation were recorded. The interviewer’s role was limited to reduce the possibility of bias.
@@ III-C Data processing
Based on the collection of documents where the transcribed conversation data was saved, we further processed the data so that it could be used for further analysis in our keyword search algorithm. We processed each transcript by applying a series of text preprocessing steps: removing newline characters, adding spaces between digits and alphabets, and combining these operations to create a cleaned version of the conversation. The text was then tokenized using the ‘$word\_tokenize$‘ function in NLTK (https://www.nltk.org) to break it into individual words. Subsequently, a text cleaning function was applied, removing short words, eliminating spurious characters, transforming letters to lowercase, and removing stopwords and specific words defined in the process, such as the names of participants. The final step involved lemmatization, which normalizes the tokens using NLTK’s WordNet lemmatizer. To illustrate, consider a sample text such as ”So changing lanes and avoiding blind spot collisions is very, very good here. I’m maintaining speeds and safe speeds.”. After applying the data processing, the lemmatized tokenized data results will be formatted as a list: [’changing’, ’lane’, ’avoiding’, ’blind’, ’spot’, ’collision’, ’good’, ’maintaining’, ’speed’, ’safe’, ’speed’].
@@ III-D Seed words and keyword search
To identify instances in the transcribed conversation data where interview participants mentioned the driving tasks and their perceived safety and trust, we employed seed words and a keyword search algorithm with the data. Seed words were defined as individual terms associated with driving tasks, perceived safety, and trust. For driving tasks, we defined the seed words as the verb corresponding to each driving task, except for steering. The choice of the seed word ”steer” was often connected to the expression ”steering wheel,” which was not our focus. Therefore, we sought alternative terms describing the steering process, namely ’change lane’ and ’keep lane’. Defining seed words is an iterative step, aligning with [25]’s suggestion to derive seeds based on the researcher’s subject knowledge. For perceived safety and trust, we adapted the seed words used by [22]. Subsequently, our keyword search algorithm, which operates by using the defined seed words as keywords, was employed to identify instances in each conversation where participants mentioned these seed words. We chose keyword search over manual inspection because relevant content could potentially span across various answers to different questions, and the considerable length of the conversations made manual inspection impractical. The seed words and the algorithm can be found in Table I and Algorithm 1. In Table I, the use of ’and’ indicates that both words must be present for the keyword search algorithm, while the use of ’or’ indicates that the presence of either one is sufficient. As an example, assume we have a list of lemmatized tokenized data: [’vehicle’, ’good’, ’brake’, ’really’, ’hard’, ’try’, ’make’, ’left’, ’turn’, ’reason’, ’turn’, ’blinker’, ’quick’, ’definitely’, ’make’, ’feel’, ’unsafe’, ’house’, ’spot’, ’signal’, ’make’, ’better’, ’before’]. We are looking for words discussing perceived safety and trust, particularly those related to braking. We set ’brake’ as the seed word for braking and ’safe,’ ’happy,’ ’relax,’ ’comfort,’ ’glad,’ ’trust,’ ’distract,’, ’rely,’ and ’trustworthy’ as seed words for perceived safety and trust. The algorithm goes through the list, searching for occurrences of these seed words. When it finds one occurrence in one of the seed words, for instance, the word ’brake’, it will expand the search over the next 20 words, resulting in a new list of words [’brake’, ’really’, ’hard’, ’try’, ’make’, ’left’, ’turn’, ’reason’, ’turn’, ’blinker’, ’quick’, ’definitely’, ’make’, ’feel’, ’unsafe’, ’house’, ’spot’, ’signal’, ’make’]. Then, the algorithm checks for any words related to the seeds of perceived safety and trust, and it identifies the word ’safe’ inside the term ’unsafe’. Now, there are two occurrences of seed words in the list, and the list will be retrieved.
After defining seed words and implementing keyword search, we qualitatively classified perceptions by analyzing the word context surrounding the keywords in the transcribed conversations. The classification process began with an initial assessment conducted by the first author. Subsequently, to minimize subjectivity in categorization, discussions were held with the co-authors to ensure alignment with others’ perspectives. These discussions involved verifying whether the systems tracked participants’ reasons during actions such as braking and determining whether the braking process indicated positive or negative perceptions of safety and trust.
@@ IV-A1 Controversy: lane changing
The participants mentioned thirteen instances where they felt safe with the systems. Nine instances indicated that the systems tracked participants’ reasons, while the rest did not. One participant emphasized that significant improvements in automatic lane changes contributed to their sense of safety. Previously, their vehicle would perform several unsafe lane changes, but now it has become as safe as their own driving. Another participant simply did not want to be bothered with lane-changing, as they believed that the systems worked properly and safely.
  • ”I do feel safe when it’s doing automatic lane changes, and that’s a massive improvement since when it first was released, it used like almost every autopilot lane change when it first came out. Would be cutting someone off. Pulling into a lane with someone rapidly approaching or some other form of unsafe lane change now..The automatic lane change feature is.. about as safe as I am. (R007)”
@@
  • ”I feel safe when I am on the freeway.. it’s s better driver than I am at taking turns on the freeway.. Changing lanes.. I like how it keeps his turn signal on all the way through the entire lane change and sometimes human drivers will just do a couple blinks. (R043)”
@@ IV-A3 Reasons mostly tracked: lane keeping
When the participants perceived lane keeping as safe, all instances showed that their reasons were tracked. Participants mentioned eleven instances where the automated driving system’s lane keeping tracked their reasons and gave them safe perceptions. One participant described that they felt safe because the systems reduced their main workload so that they did not have to pay attention to lane keeping. Another participant mentioned that the systems did a very good job of lane keeping. For example:
  • ”I feel like Autopilot makes me safer because it reduces my workload.. in terms of lane keeping mainly.. It makes me feel more comfortable as a driver just because I don’t have to pay attention to the lines on the road. (R024)”
@@ IV-B1 Reasons mostly not tracked: lane changing
When the participants explained the situations in which they felt unsafe while changing lanes, the automated driving systems did not track their reasons at all. One of the participants mentioned that they did not feel safe because the system did not have human-like behaviors. They further emphasized that it was too fast or too close to other vehicles at stop signs or traffic lights. Another participant highlighted that they could feel safer if the systems could adapt to how humans drive and incorporate a more human-like feeling into them.
@@ IV-D1 Reasons mostly not tracked: braking
Participants who felt lack of trust described that the systems failed to track their reasons. They mentioned that they did not trust the automated driving systems because the systems could make unexpected wrong decisions very fast, such as braking when there was a pedestrian on the crosswalks. Another participant described that they did not trust the systems when entering a highway because the systems did not brake when there was not much of a gap.
  • ”I don’t fully trust it when it’s active.. it can just do something wrong very fast.. you just do not expect it. For example, if it sees a pedestrian.. on the crosswalk, it’ll kind of just slam on the brakes. (R065)”
@@
  • ”I think the Autopilot may even be better than a human, and if we take emergency braking, then it can probably react and brake faster than me or certainly close to that, but.. you have to assess the situation, then it’s uncertain, and that’s why I always have to rely on myself. (R015)”
@@ V-A1 Perceived safety
On the other hand, feeling safe doesn’t always mean that the automated driving systems tracked all their reasons; there were instances of lane-changing and braking tasks where tracking failed. One participant noted that braking was too slow for them, indicating their expected deceleration reason was not tracked. However, they felt safe as long as they were ready to take over. When the driver realized operational reasons for deceleration were not fulfilled, they often took control. The participant also reported no difficulties with the takeover process. [27] found that aggressive drivers were more likely to take control when driving defensively programmed automated vehicles (AVs), potentially explaining the frequent initiation of takeover processes by drivers. The study also revealed that perceived safety levels were similar when aggressive drivers operated both aggressive and defensive AVs. Despite this, the safety score remained higher than when defensive drivers operated aggressive AVs, potentially explaining the continued feeling of safety.
The other instances indicated that the vehicle acted differently from what they expected, but it still led to perceived safety because they believed it was a better driver. The experience of many unexpected maneuvers that turned out to be correct decisions also influenced their belief. The more the vehicles behaved according to the expected reasons, the more they gained trust from the driver. This trust would make people perceive the system as safe, even if it fails to track their reasons. This is in line with the findings on the effect of reliable experiences with trust that the system is safe [26]. Thus, we found indications that tracking driver expectations, such as performing safe maneuvers like lane-keeping, braking, and lane-changing like human drivers, positively correlated with perceived safety. However, failure to track reasons for lane changing and braking was not necessarily associated with perceived lack of safety; it depended on other factors like the numerous reliable experiences, the driver’s trust level, and ease of taking over control.
@@ V-B1 Trust
We found indications that tracking driver expectations, including improved maneuver execution and human-like performance in intended traffic situations in lane-changing and braking tasks, positively relates to trust. Additionally, transparency, which does not relate to tactical or operational reasons, was also positively associated with trust. Notably, the failure to track reasons for braking tasks did not always result in the lack of trust; factors such as positive experiences and the ease of taking over control played a crucial role here.
@@ V-C Limitations
First, the data in this research focused on drivers’ subjective perceptions which may not accurately reflect the actual on-road situation (e.g., findings from [34] indicate that subjective risk can significantly differ from actual crash risk). Future research should investigate telemetry data in addition to subjective reports. Second, we limited the tactical and operational reasons for relevant human agents only to the driver’s perspective. However, many other agents (e.g., vehicle manufacturers and other road users) could and should have meaningful human control over automated driving systems [12]. We recommend further research to evaluate the tracking of the reasons of other relevant human agents for a more holistic analysis. Third, our results (e.g., Table II) might not have fully captured the entirety of the participants’ responses. The use of seed words may restrict instances that are essentially the same but articulated with different words. We recommend incorporating a more extensive set of alternative seed words, especially in the context of driving tasks, for more comprehensive results. Fourth, our findings are limited by the small number of participants in the FSD Beta program. Future research should aim for a more diverse sample encompassing a wider range of manufacturers and participants to better represent the population. Finally, the qualitative nature of our evaluation presents challenges in scalability, particularly when confronted with a larger database of interviews. Given the expansive volume of data, our current methods may prove impractical to implement.
@@ Acknowledgments
We declare that OpenAI ChatGPT was used in the writing process to enhance the style and conciseness of the text originally written by the authors.
This work was funded by the Indonesia Endowment Fund for Education (LPDP) under Grant 0006552/TRA/D/19/lpdp2021.